Six uncharged capacitors with equal capacitances are combined in parallel. The combination is connected to a 5.25-V battery, which charges the capacitors. The charging process involves 1.19 × 10-4 C of charge moving through the battery. Find the capacitance of each capacitor.
Answer #1

Let the capacitance of eachcapacitor is 'c'

when the capacitors are connected in parallel then the total capacitance is the sum of all individual capacitances

total capacitance is 6*c

Given potential difference v = 5.25 V

total charge is Q = 1.19 * 10^-4 C

We have the relation

Q = C * v

Q = 6 * c * v

1.19 * 10^-4 = 6 * c * 5.25

c = 3.77 * 10^-6 F (capacitance of each capacitor)